# On total transitivity of graphs

### v1

@@ 3 Total Transitivity in split graphs
Clearly, $Tr_{t}(G)\geq 1$. Also, from the definition of total transitivity, we know that $Tr_{t}(G)\leq Tr(G)$. Again, from [18], we know $Tr(G)\leq\omega(G)+1$. Therefore, for a split graph, $Tr_{t}(G)\leq\omega(G)+1$. Now we show that $Tr_{t}(G)$ cannot be $\omega(G)+1$ or $\omega(G)$. Let $Tr_{t}(G)=\omega(G)+1$ and $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)+1}\}$ be a total transitive partition of $G$. Since $|K|=\omega(G)$, there exists at least one set in $\pi$, say $V_{i}$, such that $V_{i}$ contains only vertices from $S$. Note that, in this case, $V_{i}$ cannot dominate $V_{i}$ as $S$ is an independent set of $G$. Therefore, we have a contradiction. Hence, $Tr_{t}(G)<\omega(G)+1$.
[…]
@@
    The $deg(s)=1$ for all $s\in S$ and $K$ has at most vertex without a neighbour in $S$.
@@
Let $G$ satisfy the condition $(a)$. We show that $Tr_{t}(G)=1$. Assume $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ is a total transitive partition of $G$. Since $deg(s)=1$ for all $s\in S$, no vertices from $S$ can be in $V_{p}$ for $p\geq 2$. To dominate vertices from $S$, $V_{1}$ contains all vertices from $K$ except at most one. So $t$ cannot be more than $2$, and $V_{2}$ can contain at most one vertex from $K$, which is a contradiction as $\pi$ is a total transitive partition. Hence, $Tr_{t}(G)=1$.
For the condition $(b)$, let $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ is a total transitive partition of $G$. Clearly, if the $deg(s)=1$ for some $s\in S$, it must belong to $V_{1}$, so assume there exists $s^{\prime}\in S$ and $deg(s^{\prime})\geq 2$. By the condition $(b)$, each vertex from $N(s^{\prime})$ is a support vertex. Therefore, each vertex from $N(s^{\prime})$ must be in $V_{1}$, and hence $s^{\prime}$ cannot be in $V_{p}$ for $p\geq 2$. From above, we have all vertices must be in $V_{1}$, and as $K$ has at most one vertex without a neighbour in $S$, $V_{p}$ must be empty from $p\geq 2$. Hence, $Tr_{t}(G)=1$.
@@
We show that $K$ has at most one vertex without a neighbour in $S$. If $K$ has more than one vertex without a neighbour in $S$, then we show a contradiction. Let $\{k_{1},k_{2}\}\subseteq K$ such that $k_{1},k_{2}$ has no neighbour in $S$. Let us consider a vertex partition $\{V_{1},V_{2}\}$ of $G$ as follows: $V_{2}=\{k_{1},k_{2}\}$, and $V_{1}$ contains all the other vertices of $G$. Clearly, $\{V_{1},V_{2}\}$ is a total transitive partition of $G$ with size $2$, which implies that $Tr_{t}(G)\geq 2$, a contradiction as $Tr_{t}(G)=1$. Hence, $K$ has at most one vertex without a neighbour in $S$, and $G$ satisfies condition $(a)$.
@@
Let $s\in S$ and $deg(s)\geq 2$. Also, let $N(s)=\{k_{1},k_{2},\ldots,k_{q}\}$, where $k_{i}\in K$. If there exists a $j$ such that $k_{j}$ is not a support vertex, then $\{V_{1},V_{2}\}$, where $V_{2}=\{s,k_{j}\}$ and $V_{1}$ contains all the other vertices of $G$, form a total transitive partition of $G$ of size $2$. Therefore, we have a contradiction as $Tr_{t}(G)=1$. So, all the vertices from $N(s)=\{k_{1},k_{2},\ldots,k_{q}\}$ are support vertex. Again, if $K$ has more than one vertex without a neighbour in $S$, similar to Case 1, we can show a contradiction. Hence, in this case, for all $s\in S$ with $deg(s)\geq 2$, each vertex from $N(s)$ is a support vertex, and $K$ has at most one vertex without a neighbour in $S$; that is, $G$ satisfies condition $(b)$. ∎
Next, we characterize split graphs with $Tr_{t}(G)=\omega(G)-1$. Before characterizing the split graphs with $Tr_{t}(G)=\omega(G)-1$, we have the following observations for any $p$, $11\leq p\leq q-$ 1, where $q=\omega(G)$.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of a split graph $G$. If $V_{j}$, $j\neq 1$ contains more than one vertices from $S$, then there exists $\pi^{\prime}=\{V_{1}^{\prime},V_{2}^{\prime},\ldots,V_{p}^{\prime}\}$ such that $|V_{j}^{\prime}\cap S|\leq 1$ for all $j\neq 1$. Moreover, when $p=\omega(G)-1$, there exists a total transitive partition of $G$, say $\pi^{\prime\prime}=\{V_{1}^{\prime\prime},V_{2}^{\prime\prime},\ldots,V_{p}^{\prime\prime}\}$ such that $|V_{i}|=2$ for all $2\leq i\leq p$.
@@
From Observation 18, we have if $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ is a total transitive partition of $G$. Then there must exist a subset of $K$, say $K^{\prime}$, such that $K^{\prime}$ dominates $S$, and all the vertices of $K^{\prime}$ must be in $V_{1}$. Moreover, $11\leq|K^{\prime}|\leq q-p+$ 1, that is, the size of the $K^{\prime}$ is at most $q-p+1$. Let us define the minimum cardinality of $K^{\prime}$ as S domination number by K and denote it by $dom_{K}(S)$ and $K^{\prime}$ with minimum cardinality as $dom_{K}(S)$-set.
Let $G$ be a split graph, and $Tr_{t}(G)=p$ for some $11\leq p\leq\omega(G)-$ 1. Also, assume that $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ with size $p$. Then we say $dom_{K}(S)$-set exists if there exists a subset $K^{\prime}$ of $K$, such that $K^{\prime}$ dominates $S$, vertices of $K^{\prime}$ are in $V_{1}$ and $11\leq|K^{\prime}|\leq q-p+$ 1, where $q=\omega(G)$.
@@
    $dom_{K}(S)$-set exists and $dom_{K}(S)=1$. Then there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$ Also, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set.
@@
    $dom_{K}(S)$-set exists and $dom_{K}(S)=2$. Then there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$ Also, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set.
@@
Let $G$ be a split graph, and $G$ satisfies the condition $(a)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(a)$, there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$ Also, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},s_{1}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, $V_{\omega(G)-1}=\{x,y\}$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1}\}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{j}\}\subseteq N(s_{j})$, we have $V_{s}$ dominates $V_{t}$ for all $22\leq s\leq t\leq\omega(G)-$ 2. Furthermore, as $V_{\omega(G)-1}=\{x,y\}$ and $K$ is a clique, $V_{s}$ dominates $V_{\omega(G)-1}$ for all $11\leq s\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma13, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(a)$, $Tr_{t}(G)=\omega(G)-1$.
Now assume $G$ be a split graph and $G$ satisfies the condition $(b)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(b)$, there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$ Also, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, we have $V_{s}$ dominates $V_{t}$ for all $12\leq s\leq t\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma13, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(b)$, $Tr_{t}(G)=\omega(G)-1$.
Therefore, if $G$ is a split graph and satisfies either $(a)$ or $(b)$ then $Tr_{t}(G)=\omega(G)-1$.
@@
By the Observation17, we can take a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1}\}\subseteq K$ be a $dom_{K}(S)$-set. Again from Observation15, we know that each $V_{i}$ must contain a vertex from $K$. Therefore, exactly one set, say $V_{r}$ from $\pi$, contains two vertices from $K$. First, assume $V_{r}$ is not $V_{1}$, then $V_{1}$ contains exactly one vertex from $K$ and at least one vertex from $S$. Also, each set $V_{2},\ldots,V_{r-1},V_{r+1},\ldots,V_{p}$ contains exactly one vertex from $K$ and exactly one vertex from $S$. Without loss of generality, assume $\{s_{1},k_{1}\}\subseteq V_{1}$ and $V_{j}=\{s_{j},k_{j}\}$ for all all $1\leq j\leq r-1,r+1\leq j\leq p$ and $V_{r}=\{k_{r},k_{r}^{\prime}\}$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{j})\geq j$ for all $1\leq j\leq p$ and $j\neq r$. Moreover, $\{k_{1},k_{2},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $11\leq j\leq r-$ 1 and $\{k_{1},k_{2},\ldots,k_{r-1},k_{r+1},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $r+1\leq j\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{r},k_{r}^{\prime}\}$. Similarly, we can prove that $G$ satisfies the condition $(a)$ when $V_{r}$ is $V_{1}$.
@@
In this case, we can also assume there exists a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1},k_{1}^{\prime}\}\subseteq K$ be a $dom_{K}(S)$-set. From Observation15 we know that each $V_{i}$ must contains a vertex from $K$ and by definition of $dom_{K}(S)$-set $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$. So, each $V_{i}$ contains exactly one vertex from $K$ and exactly one vertex from $S$, for all $2\leq i\leq p$. Without loss of generality assume $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ and $V_{i}=\{s_{i},k_{i}\}$ for all all $1\leq i\leq p$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{i})\geq j$ for all $2\leq i\leq p$. Moreover, $\{k_{2},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $1\leq i\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$.
@@
Let $Tr_{t}(G)=p$ and $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ with size $p$. By the Observation18, $dom_{K}(S)$-set exists and $11\leq|dom_{K}(S)|\leq p-q+$ 1. We divide our proof into two cases:
@@
As $\pi$ is a total transitive partition, size each $|V_{i}|\geq 2$. Also, by Observation15, each $V_{i}$ must contain a vertex from $K$. Now it is given that $Tr_{t}(G)=p$ and $|K|=q$, so each $p-(q-p)=2p-q$ set from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq p-(q-p)=2p-q$ $(a(i))$. Again, by the Observation16, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\alpha}}\}\subseteq S$, where $\alpha=2p-q$ such that $s_{i_{j}}\in V_{i_{j}}$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(a(ii))$.

### v2

@@ 3 Total Transitivity in split graphs
Clearly, $Tr_{t}(G)\geq 1$. Also, from the definition of total transitivity, we know that $Tr_{t}(G)\leq Tr(G)$. Again, from [19], we know $Tr(G)\leq\omega(G)+1$. Therefore, for a split graph, $Tr_{t}(G)\leq\omega(G)+1$. Now we show that $Tr_{t}(G)$ cannot be $\omega(G)+1$ or $\omega(G)$. Let $Tr_{t}(G)=\omega(G)+1$ and $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)+1}\}$ be a total transitive partition of $G$. Since $|K|=\omega(G)$, there exists at least one set in $\pi$, say $V_{i}$, such that $V_{i}$ contains only vertices from $S$. Note that, in this case, $V_{i}$ cannot totally dominate $V_{i}$ as $S$ is an independent set of $G$. Therefore, we have a contradiction. Hence, $Tr_{t}(G)<\omega(G)+1$.
Further assume $Tr_{t}(G)=\omega(G)$ and $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)}\}$ is a total transitive partition of $G$. By the Proposition 9, we can assume that $|V_{\omega(G)}|=2$. Now $V_{\omega(G)}$ contains either two vertices from $K$ or one vertex from $K$ and another vertex from $S$. For the first case, there exists at least one set in $\pi$, say $V_{i}$, such that $V_{i}$ contains only vertices from $S$. Because $S$ is independent, $V_{i}$ cannot totally dominate $V_{i}$, so we have a contradiction. Consider the other case, namely that $V_{\omega(G)}$ contains one vertex from $K$ and another vertex from $S$. Let $s\in S\cap V_{\omega(G)}$. Then, by the definition of total transitive partition, $deg(s)\geq\omega(G)$. But $|K|=\omega(G)$ implies that $deg(s)<\omega(G)$, which is a contradiction. Therefore, in each case, we have a contradiction. Hence, $11\leq Tr_{t}(G)\leq\omega(G)-$ 1. ∎
@@
    For all $s\in S$, $deg(s)=1$, and $K$ has at most one vertex without a neighbour in $S$.
@@
Let G satisfy the condition (a). We show that $Tr_{t}(G)=1$. Assume $\pi=\{V_{1},V_{2},\ldots,V_{t}\}$ is a total transitive partition of $G$. Since $deg(s)=1$ for all $s\in S$, no vertices from $S$ can be in $V_{p}$ for $p\geq 2$. To totally dominate vertices from $S$, $V_{1}$ contains all vertices from $K$ except at most one. Thus $V_{2}$ can contain at most one vertex from $K$, which contradicts the fact that $\pi$ is a total transitive partition. Therefore, $t$ cannot be more than $2$. Hence, $Tr_{t}(G)=1$.
For condition $(b)$, let $\pi=\{V_{1},V_{2},\dots,V_{t}\}$ be a total transitive partition of $G$. Clearly, if $deg(s)=1$ for some $s\in S$, then $s$ must belong to $V_{1}$. Now, assume there exists $s^{\prime}\in S$ such that $deg(s^{\prime})\geq 2$. By condition $(b)$, each vertex in $N(s^{\prime})$ is a support vertex. Therefore, each vertex in $N(s^{\prime})$ must be in $V_{1}$, and hence $s^{\prime}$ cannot be in $V_{p}$ for $p\geq 2$. Thus, every vertex in $S$ must be in $V_{1}$, and at most one vertex from $K$ can be in $V_{1}$. This contradicts the fact that $\pi$ is a total transitive partition. Therefore, $t$ cannot be greater than 2. Hence, $Tr_{t}(G)=1$.
@@
We show that $K$ has at most one vertex without a neighbour in $S$. If $K$ has more than one vertex without a neighbour in $S$, we will derive a contradiction. Let $\{k_{1},k_{2}\}\subseteq K$ such that $k_{1}$ and $k_{2}$ have no neighbour in $S$. Consider a vertex partition $\{V_{1},V_{2}\}$ of $G$ as follows: $V_{2}=\{k_{1},k_{2}\}$, and $V_{1}$ contains all the other vertices of $G$. Clearly, $\{V_{1},V_{2}\}$ is a total transitive partition of $G$ with size 2, which implies that $Tr_{t}(G)\geq 2$, a contradiction since $Tr_{t}(G)=1$. Hence, $K$ has at most one vertex without a neighbour in $S$, and $G$ satisfies condition $(a)$.
@@
Let $s\in S$ and $\deg(s)\geq 2$. Furthermore, let $N(s)=\{k_{1},k_{2},\dots,k_{q}\}$, where $k_{i}\in K$. If there exists a $j$ such that $k_{j}$ is not a support vertex, then $\{V_{1},V_{2}\}$, where $V_{2}=\{s,k_{j}\}$ and $V_{1}$ contains all the other vertices of $G$, forms a total transitive partition of $G$ of size 2. Therefore, we have a contradiction since $Tr_{t}(G)=1$. Thus, all the vertices from $N(s)=\{k_{1},k_{2},\dots,k_{q}\}$ must be support vertices. Similarly, if $K$ has more than one vertex without a neighbour in $S$, as in Case 1, we can show a contradiction. Hence, for all $s\in S$ with $\deg(s)\geq 2$, each vertex in $N(s)$ is a support vertex, and $K$ has at most one vertex without a neighbour in $S$. This implies that $G$ satisfies condition $(b)$. ∎
Next, we characterize split graphs with $Tr_{t}(G)=\omega(G)-1$. Before characterizing split graphs with $Tr_{t}(G)=\omega(G)-1$, we have following observations for any $p$, $11\leq p\leq q-$ 1, where $q=\omega(G)$.
Observation 14.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of a split graph $G$. If $V_{j}$, $j\neq 1$ contains more than one vertex from $S$, then there exists $\pi^{\prime}=\{V_{1}^{\prime},V_{2}^{\prime},\ldots,V_{p}^{\prime}\}$ such that $|V_{j}^{\prime}\cap S|\leq 1$ for all $j\neq 1$. Moreover, when $p=\omega(G)-1$, there exists a total transitive partition of $G$, say $\pi^{\prime\prime}=\{V_{1}^{\prime\prime},V_{2}^{\prime\prime},\ldots,V_{p}^{\prime\prime}\}$ such that $|V_{i}|=2$ for all $2\leq i\leq p$.
@@
From Observation 17, we have if $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ is a total transitive partition of $G$. Then there must exist a subset of $K$, say $K^{\prime}$, such that $K^{\prime}$ dominates $S$, and all the vertices of $K^{\prime}$ must be in $V_{1}$. Moreover, $11\leq|K^{\prime}|\leq q-p+$ 1, that is, the size of the $K^{\prime}$ is at most $q-p+1$. Let us define the minimum cardinality of $K^{\prime}$ as S domination number by K and denote it by $dom_{K}(S)$ and $K^{\prime}$ with minimum cardinality as $dom_{K}(S)$-set.
Let $G$ be a split graph, and $Tr_{t}(G)=p$ for some $11\leq p\leq\omega(G)-$ 1. Also, assume that $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ is a total transitive partition of $G$ with size $p$. Then we say $dom_{K}(S)$-set exists if there exists a subset $K^{\prime}$ of $K$, such that $K^{\prime}$ dominates $S$, vertices of $K^{\prime}$ are in $V_{1}$ and $11\leq|K^{\prime}|\leq q-p+$ 1, where $q=\omega(G)$.
@@
    $dom_{K}(S)$-set exists and $|dom_{K}(S)|=1$. Also, there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$ Furthermore, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set.
@@
    $dom_{K}(S)$-set exists and $|dom_{K}(S)|=2$. Also, there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$ Furthermore, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set.
@@
Let $G$ be a split graph, and let $G$ satisfy condition $(a)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(a)$, there exist $\{s_{1},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $1\leq j\leq\alpha$, where $\alpha=\omega(G)-2$. Furthermore, for some pair of vertices $x,y$, we have an ordering of the vertices of $K\setminus\{k_{1},x,y\}$, say $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},s_{1}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, $V_{\omega(G)-1}=\{x,y\}$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1}\}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ totally dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{j}\}\subseteq N(s_{j})$, we have $V_{s}$ totally dominates $V_{t}$ for all $22\leq s\leq t\leq\omega(G)-$ 2. Furthermore, as $V_{\omega(G)-1}=\{x,y\}$ and $K$ is a clique, $V_{s}$ totally dominates $V_{\omega(G)-1}$ for all $11\leq s\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma 12, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(a)$, $Tr_{t}(G)=\omega(G)-1$.
Now assume $G$ be a split graph and $G$ satisfies the condition $(b)$. We show that $Tr_{t}(G)=\omega(G)-1$. From the condition $(b)$, there exist $\{s_{2},\ldots,s_{\alpha}\}\subseteq S$ such that $deg(s_{j})\geq j$ for all $2\leq j\leq\alpha$, where $\alpha=\omega(G)-1$. Also, we have an ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$, say, $\{k_{2},k_{3},\ldots,k_{\alpha}\}$ such that $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, where $\{k_{1},k_{1}^{\prime}\}$ is a $dom_{K}(S)$-set. Let us consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{\omega(G)-1}\}$ of $G$ as follows: $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$, $V_{j}=\{s_{j},k_{j}\}$, for all $2\leq j\leq\alpha$, and all the other vertices of $G$ in $V_{1}$. We show that $\pi$ is a total transitive partition. Since $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ is a $dom_{K}(S)$-set and $K$ is a clique, $V_{1}$ totally dominates $V_{i}$ for all $i\geq 1$. Again, from condition $\{k_{2},k_{3},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $2\leq j\leq\alpha$, we have $V_{s}$ totally dominates $V_{t}$ for all $12\leq s\leq t\leq\omega(G)-$ 1. Therefore, $\pi$ is a total transitive partition of size $\omega(G)-1$, which implies $Tr_{t}(G)\geq\omega(G)-1$. Moreover, from Lemma 12, we have $Tr_{t}(G)\leq\omega(G)-1$. Hence, if $G$ satisfies condition $(b)$, $Tr_{t}(G)=\omega(G)-1$. Therefore, if $G$ is a split graph and satisfies either $(a)$ or $(b)$ then $Tr_{t}(G)=\omega(G)-1$.
@@
By the Observation 15, we can take a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1}\}\subseteq K$ be a $dom_{K}(S)$-set. Again from Observation 14, we know that each $V_{i}$ must contain a vertex from $K$. Therefore, exactly one set, say $V_{r}$ from $\pi$, contains two vertices from $K$. First, assume $V_{r}$ is not $V_{1}$, then $V_{1}$ contains exactly one vertex from $K$ and at least one vertex from $S$. Also, each set $V_{2},\ldots,V_{r-1},V_{r+1},\ldots,V_{p}$ contains exactly one vertex from $K$ and exactly one vertex from $S$. Without loss of generality, assume $\{s_{1},k_{1}\}\subseteq V_{1}$ and $V_{j}=\{s_{j},k_{j}\}$ for all all $2\leq j\leq r-1,r+1\leq j\leq p$ and $V_{r}=\{k_{r},k_{r}^{\prime}\}$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{j})\geq j$ for all $1\leq j\leq p$ and $j\neq r$. Moreover, $\{k_{1},k_{2},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $11\leq j\leq r-$ 1 and $\{k_{1},k_{2},\ldots,k_{r-1},k_{r+1},\ldots,k_{j}\}\subseteq N(s_{j})$ for all $r+1\leq j\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{r},k_{r}^{\prime}\}$. Similarly, we can prove that $G$ satisfies the condition $(a)$ when $V_{r}$ is $V_{1}$.
@@
In this case, we can also assume there exists a total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ of $G$ such that $|V_{i}|=2$ for all $2\leq i\leq p$, where $p=\omega(G)-1$. Let $\{k_{1},k_{1}^{\prime}\}\subseteq K$ be a $dom_{K}(S)$-set. From Observation 14 we know that each $V_{i}$ must contains a vertex from $K$ and by definition of $dom_{K}(S)$-set $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$. So, each $V_{i}$ contains exactly one vertex from $K$ and exactly one vertex from $S$, for all $2\leq i\leq p$. Without loss of generality assume $\{k_{1},k_{1}^{\prime}\}\subseteq V_{1}$ and $V_{i}=\{s_{i},k_{i}\}$ for all all $1\leq i\leq p$. Since $\pi$ is a total transitive partition of $G$ the $deg(s_{i})\geq j$ for all $2\leq i\leq p$. Moreover, $\{k_{2},\ldots,k_{i}\}\subseteq N(s_{i})$ for all $1\leq i\leq p$. Clearly, we have the required ordering of the vertices of $K\setminus\{k_{1},k_{1}^{\prime}\}$.
@@
Let $Tr_{t}(G)=p$ and $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ be a total transitive partition of $G$ with size $p$. By the Observation 17, $dom_{K}(S)$-set exists and $11\leq|dom_{K}(S)|\leq p-q+$ 1. We divide our proof into two cases:
@@
As $\pi$ is a total transitive partition, each $|V_{i}|\geq 2$. Also, by Observation 14, each $V_{i}$ must contain a vertex from $K$. Now it is given that $Tr_{t}(G)=p$ and $|K|=q$, so each $p-(q-p)=2p-q$ sets from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq p-(q-p)=2p-q$ $(a(i))$. Again, by the Observation 15, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\alpha}}\}\subseteq S$, where $\alpha=2p-q$ such that $s_{i_{j}}\in V_{i_{j}}$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(a(ii))$.